OpenStudy (anonymous):

How do you integrate the integral of (x^3) * (sin(x^2)) dx

OpenStudy (anonymous):

\[\int x^3 \sin{(x^2)}dx\] Have you tried integration by parts?

OpenStudy (anonymous):

\(u = x^2\)

OpenStudy (anonymous):

\(du = 2xdx\)

OpenStudy (anonymous):

i would write it as \[\int x^2\times x\sin(x^2)dx\] and then try parts

OpenStudy (anonymous):

what @wio said you get \[\frac{1}{2}\int u \sin(u)du\] parts for this one

OpenStudy (anonymous):

Thank you both for your help!